OpenStudy (anonymous):

f and g are defined for all real numbers x as follows. f(x)=x-6 g(x)=3x^2 write the expressions for (f-g)(x) and (f+g)(x) and evaluate (f times g)(3) -I need help understanding the steps to these questions... my teacher didn't do a good job whatsoever on explaining..

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

f(x) - g(x) = -3x^2 - x - 6 f(x) + g(x) = 3x^2 - x - 6 f(3) *g(3) = (3-6)(33^2) = -3*3*9= -81